Please help! Just part B! Thanks!Bob, whose mass is 90 kg, is out in a 4 meter long
Answer #1

let the left end be at a distance x=d from the shore
then right end will be at x=d+4

and the center of the boat will be at x=d+2

now bob is at x=d+4

hence center of mass=[90*(d+4)+110*(d+2)]/(90+110)=(200*d+580)/200


as there is no external force on the system, the center of mass is not changed.

hence (200*d+580)/200=11.1

d=8.2 m

hene the left end of the boat is at a distance of 8.2 m from the shore.